### v2

# On the Diameter of Arrangements of Topological Disks

Abstract
Let $\mathcal{D}=\{D_{0},\ldots,D_{n-1}\}$ be a set of $n$ topological disks in the plane and let $\mathcal{A}:=\mathcal{A}(\mathcal{D})$ be the arrangement induced by $\mathcal{D}$. For two disks $D_{i},D_{j}\in\mathcal{D}$, let $\Delta_{ij}$ be the number of connected components of $D_{i}\cap D_{j}$, and let $\Delta:=\max_{i,j}\Delta_{ij}$. We show that the diameter of $\mathcal{G}^{*}$, the dual graph of $\mathcal{A}$, can be bounded as a function of $n$ and $\Delta$. Thus, any two points in the plane can be connected by a Jordan curve that crosses the disk boundaries a number of times bounded by a function of $n$ and $\Delta$. In particular, for the case of two disks, we prove that the diameter of $\mathcal{G}^{*}$ is at most $\max\{2,2\Delta\}$ and this bound is tight. For the general case of $n>2$ disks, we show that the diameter of $\mathcal{G}^{*}$ is $O(n^{3}2^{n}\Delta)$. We achieve this by proving that the number of maximal faces in $\mathcal{A}$—faces whose ply is more than the ply of their neighboring faces—is $O(n^{2}2^{n}\Delta)$. To this end, we first show that the number of maximum faces—faces whose ply is nn—is $O(n^{2}\Delta)$; the latter bound, which is of independent interest, is tight in the worst case.
keywords:
Topological disk, arrangement, face, Jordan curve, crossing
1 Introduction
The arrangement $\mathcal{A}(S)$ defined by a set $S$ of objects in the plane is the subdivision of the plane into faces, edges, and vertices induced by (the boundaries of) the objects. Arrangements in the plane and in higher dimensions play a fundamental role in computational geometry, since many geometric problems can be transformed into problems on arrangements. As a result, the combinatorial complexity of arrangements—the total number of vertices, edges, and faces (and higher-dimensional features)—induced by various types of objects has been studied extensively. These studies concern the complexity of complete arrangements as well as the complexity of various substructures: single cells, many cells, zones, levels, and so on; see the book by Sharir and Agarwal [DBLP:books/daglib/0080837] or the surveys by Agarwal and Sharir [DBLP:books/el/00/AgarwalS00a] or Halperin and Sharir [DBLP:reference/cg/arrangements]. The known results on arrangements typically deal with constant-complexity objects such as lines, circles, or other objects whose boundaries intersect a bounded number of times. This is not surprising, as the complexity of the arrangement induced by some set of objects obviously cannot be bounded if we do not put a bound on the number of intersections between the boundaries of any two objects in the set. In this paper we are interested in whether anything meaningful can be said about complexity of an arrangement if the number of boundary intersections of any two objects in the arrangement is not necessarily bounded.
Let $\mathcal{D}=\{D_{0},\ldots,D_{n-1}\}$ be a set of $n$ topological disks in $\mathbb{R}^{2}$. To simplify the presentation we will refer to the objects in $\mathcal{D}$ as disks, but the reader should keep on mind that the boundaries of the disks in $\mathcal{D}$ are arbitrary closed Jordan curves. In particular, the boundaries $\partial D_{i}$ and $\partial D_{j}$ of two disks $D_{i},D_{j}\in\mathcal{D}$ can intersect arbitrarily many times. We will assume that the disks in $\mathcal{D}$ are in general position. More precisely, we assume that $\partial D_{i}\cap\partial D_{j}$ consists of finitely many points, that two boundaries properly cross every time they intersect, i.e., there are no tangencies, and that no three disk boundaries have a point in common.
Let $\mathcal{A}:=\mathcal{A}(\mathcal{D})$ be the arrangement induced by $\mathcal{D}$. Let $\Delta_{ij}$ denote the number of connected components of the intersection $D_{i}\cap D_{j}$ of a pair of objects $D_{i},D_{j}\in\mathcal{D}$, and define $\Delta:=\max_{i,j}\Delta_{ij}$. We refer to $\Delta$ as the overlap number of $\mathcal{D}$. Note that the complexity of $\mathcal{A}$ cannot be bounded as a function of $n$ and $\Delta$. Indeed, even for $n=2$ and $\Delta=1$, the number of faces in $\mathcal{A}$ can be arbitrarily large, as illustrated in Figure 1(i).
Figure 1: (i) A set of two topological disks with $\Delta=1$ can induce an arrangement with arbitrarily many faces. (ii) A set of three topological disks and the dual graph of their arrangement. The arrangement has four maximal faces, whose dual nodes are shown in black. The edges incident to the node corresponding to the unbounded face are only drawn partially, to avoid cluttering the figure. If the blue disk is $D_{0}$, the red disk is $D_{1}$, and the green disk is $D_{2}$, then $\Delta_{01}=1$ and $\Delta_{02}=\Delta_{12}=3$, so $\Delta=3$.
But are there perhaps other quantities describing how complex the arrangement $\mathcal{A}$ can be, yet be bounded in terms of $n$ and $\Delta$?
The first quantity we will study is the diameter of the dual graph of the arrangement. The dual graph of the arrangement $\mathcal{A}$ is the graph $\mathcal{G}^{*}:=\mathcal{G}^{*}(\mathcal{A})$ whose nodes correspond to the faces in $\mathcal{A}$ and two nodes in $\mathcal{G}^{*}$ are adjacent if and only if the corresponding faces have at least one edge in common; see Fig. 1(ii). The diameter of $\mathcal{G}^{*}$ is the maximum hop-distance between any two nodes in $\mathcal{G}^{*}$.
The second quantity of interest is the number of maximal faces in $\mathcal{A}$. Let ply of a face in $\mathcal{A}$ be the number of disks in $\mathcal{D}$ that contain the face. Two faces are called neighbors in $\mathcal{A}$ if they have at least one common edge, if their corresponding nodes in $\mathcal{G}^{*}$ are adjacent. A face is called maximal if its ply is larger than the ply of any of its neighboring faces. A special case of a maximal face is a maximum face, which is a face whose ply is $n$ or, in other words, a face that is contained in all disks in $\mathcal{D}$.
Our results.
Let $\operatorname{diam}(\mathcal{G})$ denote the diameter of a graph $\mathcal{G}$, and let $\mu(\mathcal{A})$ and $M(\mathcal{A})$ denote the number of maximal and maximum faces, respectively, of the arrangement $\mathcal{A}$. Furthermore, define
$\xi(n,\Delta)$ $\coloneq\max_{\mathcal{D}}\{\operatorname{diam}(\mathcal{G}^{*}(\mathcal{A}(\mathcal{D}))\}$
and
$\mu(n,\Delta)$ $\coloneq\max_{\mathcal{D}}\{\mu(\mathcal{A}(\mathcal{D}))\}$
and
$M(n,\Delta)$ $\coloneq\max_{\mathcal{D}}\{M(\mathcal{A}(\mathcal{D}))\},$
with the maximum taken over all sets $\mathcal{D}$ of $n$ topological disks with overlap number $\Delta$ in general position. We first consider the case of two disks. Bounding $\mu(\mathcal{A})$ is trivial for $n=2$, since $\mu(\mathcal{A})=\Delta$ if $D_{0}$ and $D_{1}$ intersect, and $\mu(\mathcal{A})=2$ if $D_{0}$ and $D_{1}$ are disjoint. Thus, $\mu(2,\Delta)=\max\{2,\Delta\}$ and $M(2,\Delta)=\Delta$. Bounding $\xi(2,\Delta)$ is less straightforward. We study this problem in Section 2, where we prove the following tight bound on the maximum possible diameter of $\mathcal{G}^{*}(\mathcal{A}(\mathcal{D}))$ in terms of $\Delta$.
Theorem 1.1.
$\xi(2,\Delta)=\max\{2,2\Delta\}$.
In Section 3 we then turn our attention to the much more challenging case of general $n$. Here, we first prove the following upper bound on the number of maximum faces.
Theorem 1.2.
For any $n>2$, we have $M(n,\Delta)<2n(n-1)\Delta$.
We also give a construction showing that $M(n,\Delta)=\Omega(n^{2}\Delta)$, showing that the bound from Theorem 1.2 is asymptotically tight. This result implies the following bound on the number of maximal faces.
Theorem 1.3.
For any $n>2$, we have $\mu(n,\Delta)=O(n^{2}2^{n}\Delta)$.
Finally, we show that the diameter of $\mathcal{G}^{*}$ is bounded in terms of $n$ and $\Delta$.
Theorem 1.4.
For any $n>2$, we have $\xi(n,\Delta)=O(n^{3}2^{n}\Delta)$.
Related work.
The problem of bounding the diameter of $\mathcal{G}^{*}$ is equivalent to the following question. For two points $s,t\in\mathbb{R}^{2}$ not lying on the boundary of any disk in $\mathcal{D}$, let $d(s,t)$ be the minimum number of times one needs to cross a disk boundary when going from $s$ to $t$. Then, the question is: what is the maximum value of $d(s,t)$ over all such pairs $s$ and $t$? This is related to the concept of barrier coverage in sensor networks, which aims to ensure that all possible paths through a surveillance domain will intersect the boundary of a sensor region at least once. Cardei and Wu [cardei2004coverage] and Meguerdichian et al. [meguerdichian2001coverage] provide comprehensive overviews of work on this topic. The notion of barrier coverage closest to our work is the one introduced by Bereg and Kirkpatrick [bereg2009approximating], who define the thickness of a barrier as the minimum number of sensor region intersections of a path between any two regions. A substantial difference between the paper by Bereg and Kirkpatrick and our paper is the way of counting the number of crossings of an $st$-path with the region boundaries: they count a boundary crossing only if the curve enters a region for the first time, while we count every boundary crossing. Cabello et al. [DBLP:journals/ijcga/AltCGK17] study the related problem of computing, for a given set of line segments and two points $s$ and $t$, the minimum number of segments crossed by an $st$-path, but they also count each segment at most once.
Notation.
In the following, $\mathcal{D}=\{D_{0},\ldots,D_{n-1}\}$ always refers to a set of $n$ topological disks in general position. Furthermore, $\mathcal{A}$ refers to the arrangement induced by $\mathcal{D}$, and $\mathcal{G}^{*}$ to the dual graph of $\mathcal{A}$. For two disks $D_{i},D_{j}\in\mathcal{D}$, $\mathcal{C}(D_{i},D_{j})$ denotes the set of connected components of $D_{i}\cap D_{j}$. Recall that $\Delta_{ij}:=|\mathcal{C}(D_{i},D_{j})|$ denotes the number of such connected components, and that the overlap number of $\mathcal{D}$ is defined as $\Delta:=\max_{ij}\Delta_{ij}$. For two nodes $u$ and $v$ in the dual graph $\mathcal{G}^{*}$ of the arrangement $\mathcal{A}$, we use $d(u,v)$ to denote the distance between $u$ and $v$ in $\mathcal{G}^{*}$, that is, the minimum number of edges on any path from $u$ to $v$ in $\mathcal{G}^{*}$. When we speak of the hop-distance between two faces $f,f^{\prime}$ in the arrangement $\mathcal{A}$, we refer to the distance between the nodes in $\mathcal{G}^{*}$ that correspond to $f$ and $f^{\prime}$. With a slight abuse of notation, we denote this hop-distance by $d(f,f^{\prime})$.
2 The case of two disks
In this section we will prove Theorem 1.1. We first prove the upper bound — that is, we show that the diameter of the dual graph $\mathcal{G}^{*}$ of an arrangement of two topological disks is at most $\max\{2,2\Delta\}$ — and then, we provide an example showing that this bound is tight.
The upper bound.
It is easy to verify that $\operatorname{diam}(\mathcal{G}^{*})=2$ when $\Delta=0$ or $\Delta=1$, so from now on we assume that $\Delta\geqslant 2$. Let $D_{0}$ and $D_{1}$ be the two disks, and imagine that $D_{0}$ is colored red and $D_{1}$ is colored blue. We color the faces of $\mathcal{A}(\mathcal{D})$ accordingly: faces contained in $D_{0}\setminus D_{1}$ are red, faces in $D_{1}\setminus D_{0}$ are blue, faces in $D_{0}\cap D_{1}$ are purple, and faces in $\mathbb{R}^{2}\setminus(D_{0}\cup D_{1})$ are white; see Fig. 2. Similarly, we color the nodes of $\mathcal{G}^{*}$ in the colors of their respective faces. Observe that any edge in $\mathcal{G}^{*}$ is either between a white node and a red or blue node, or between a purple node and a red or blue node; there are neither edges between white and purple nodes, nor edges between red and blue nodes. Also observe that the number of purple nodes is, by definition, exactly $\Delta$. We call a path in $\mathcal{G}^{*}$ a colored path if it only uses non-white nodes.
Lemma 2.1.
For any two purple nodes $u,v$ in $\mathcal{G}^{*}$ there is a colored path from $u$ to $v$ of length at most $2\Delta-2$.
Proof 2.2.
First, notice that a colored path from $u$ to $v$ always exists since $u$ and $v$ are purple. Indeed, the corresponding faces $f_{u}$ and $f_{v}$ lie in $D_{0}\cap D_{1}$, so there is a curve $\gamma\subset D_{0}$ connecting $f_{u}$ and $f_{v}$ inside $D_{0}$. This curve corresponds to a colored path in $\mathcal{G}^{*}$.
Let $\pi(u,v)$ be a shortest colored path from $u$ to $v$ in $\mathcal{G}^{*}$. Since $\pi(u,v)$ is a shortest path, any node occurs at most once on it. Moreover, $\pi(u,v)$ does not have white nodes, so it alternates between purple nodes and red or blue nodes. Since $\pi(u,v)$ starts and ends at a purple node and it contains at most $\Delta$ purple nodes, $\pi(u,v)$ has at most $2\Delta-2$ edges.
Figure 2: The dual graph of an arrangement of two disks, with white, blue, red, and purple nodes.
Since we assumed $\Delta>0$, any red or blue node has a purple neighbor. By Lemma 2.1 the distance between any two purple nodes is at most $2\Delta-2$ that implies $d(u,v)\leqslant 2\Delta$ for any two non-white nodes $u$ and $v$. Thus, to prove that $\operatorname{diam}(\mathcal{G}^{*})\leqslant 2\Delta$, it remains to consider the case where at least one of the nodes $u,v$ is white. Observe that any white node has a red neighbor (and a blue neighbor as well), and so, it can reach a purple node in two steps. This implies that $d(u,v)\leqslant 2\Delta+2$ for any two nodes $u,v$: we can reach a purple node $w_{1}$ from $u$ in at most two steps, we can reach a purple node $w_{2}$ from $v$ in at most two steps, and $d(w_{1},w_{2})\leqslant 2\Delta-2$ by Lemma 2.1. Next we show that $d(u,v)\leqslant 2\Delta$ even if one or both of the nodes $u,v$ are white. The improvement is based on the following lemma.
Lemma 2.3.
Let $\Delta\geqslant 2$ and let $u$ be a white node in $\mathcal{G}^{*}$. Then, there are two distinct purple nodes $w,w^{\prime}$ with $d(u,w)=d(u,w^{\prime})=2$.
Proof 2.4.
Let $u$ be a white node in $\mathcal{G}^{*}$ and $f_{u}$ its corresponding face in $\mathcal{A}$. There are two cases to consider.
Case I: $f_{u}$ is a bounded face. Let $v_{1},\ldots,v_{t}$ be the vertices of $\mathcal{A}$ incident to $f_{u}$, enumerated in order along $\partial f_{u}$. Suppose for a contradiction that all purple faces incident to the vertices of $f_{u}$ are one and the same purple face $f$. Note that $d(f,f_{u})=2$. Let $x$ be a point in the interior of $f$. Draw for each vertex $v_{i}$ a curve $\gamma_{i}\subset f$ from $x$ to $v_{i}$; see Fig. 3(i). It will be convenient to draw the curves such that they are pairwise disjoint except at the shared endpoint $x$.
Figure 3: (i) The edge $e_{2}=v_{2}v_{3}$ and the curves $\gamma_{2}$ and $\gamma_{3}$ form a closed curve enclosing $f_{u}$. (ii) If the purple faces $f_{1}$ and $f_{2}$ are the same face, there is a curve $\gamma\subset f_{1}$ connecting the endpoints of $e^{*}$.
Then, there must be an edge $e_{i}=v_{i}v_{i+1}$, with indices taken modulo $t$, that, together with the two curves $\gamma_{i},\gamma_{i+1}$, forms a closed curve $\Gamma$ enclosing $f_{u}$. To see this, imagine adding the curves $\gamma_{1},\ldots,\gamma_{t}$ one by one and consider the arrangement defined by $\partial f$ (which is a simple closed curve) and the added curves $\gamma_{i}$. It is easy to argue by induction that the unbounded face of this arrangement contains an edge $e\subset\partial f_{u}$. This edge is the edge $e_{i}$ with the required properties. If $e_{i}\subset\partial D_{0}$ then $\Gamma\subset D_{0}$, and if $e_{i}\subset\partial D_{1}$ then $\Gamma\subset D_{1}$, both contradicting that $f_{u}$ is a white face.
Case II: $f_{u}$ is the unbounded face. In this case, we can assume there is a purple face $f$ with $d(f,f_{u})>2$, otherwise we are done. Pick any edge $e\subset\partial D_{0}$ of $f$. Note that $e$ is not an edge of $f_{u}$, since $f$ is purple and $f_{u}$ is white. Now consider $\left(\mathbb{R}^{2}\setminus f_{u}\right)\setminus D_{0}$. This region consists of several components, one for each edge of $f_{u}$ contributed by $\partial D_{1}$. Note that $e$ is contained in the boundary of one of these components, which we denote by $C$. Let $e^{*}:=\partial C\cap\partial D_{1}$; see Fig. 3(ii). Let $f_{1}$ and $f_{2}$ be the two purple faces incident to the endpoints of $e^{*}$. Then $f_{1}$ and $f_{2}$ must be distinct purple faces, thus proving the claim. Indeed, if $f_{1}=f_{2}$ then there is a curve $\gamma\subset f_{1}$ connecting the endpoints of $e^{*}$ such that $\gamma\cup e^{*}$ is a closed curve contained in $D_{1}$ and enclosing $e$, contradicting that $f\neq f_{1},f_{2}$, while we must have $f\neq f_{1},f_{2}$ since $d(f,f_{u})>2$.
With Lemmas 2.1 and 2.3 available, we now argue that $d(u,v)\leqslant 2\Delta$ for any two nodes $u,v$ in $\mathcal{G}^{*}$. We already proved this when both $u$ and $v$ are non-white. Now assume that $u$ is a white node. Let $d_{\max}$ be the maximum distance between any two purple nodes. We have two cases.
  • If $d_{\max}<2\Delta-2$, then $d_{\max}\leqslant 2\Delta-4$. Indeed, any path in $\mathcal{G}^{*}$ alternates between a purple or white node and a blue or red node. So, any path between two purple nodes has even length. Since both $u$ and $v$ can reach a purple node in at most two steps, we have $d(u,v)\leqslant d_{\max}+2+2\leqslant 2\Delta$.
  • If $d_{\max}=2\Delta-2$, by Lemma 2.1 there are two purple nodes $w_{1},w_{2}$ whose shortest colored path $\pi(w_{1},w_{2})$ has length $2\Delta-2$ and passes through all the purple nodes. By Lemma 2.3, in two steps $u$ can reach at least two distinct purple nodes on this path. This implies that it can reach every purple node in at most $2\Delta-2$ steps. Since $v$ can reach a purple node in at most two steps, $d(u,v)\leqslant(2\Delta-2)+2=2\Delta$.
This finishes the proof of the upper bound of Theorem 1.1.
The lower bound.
Figure 4: Left: Two spiral-shaped topological disks with overlap number $\Delta=6$. The path shown in the figure is a shortest path in the dual graph between the nodes corresponding to the innermost blue face and the outermost blue face. Its length is $2\Delta=12$. Right: Any path between the innermost and the outermost blue face must cross each of the black and gray rings.
The dual graph $\mathcal{G}^{*}$ of two disjoint disks has diameter 2. Therefore, to prove the lower bound, it suffices to construct an arrangement of two disks with the overlap number $\Delta\geqslant 2$ whose dual graph has diameter $2\Delta$. Fig. 4 exhibits such an arrangement for $\Delta=6$. Each spiral consists of $4k$ thin corridors, two vertical and two horizontal. In the figure $k=3$, but the example easily extends to larger $k$. Note that when increasing $k$ by 1, we increase $\Delta$ by 2 as each vertical corridor of the blue disk intersects a horizontal corridor of the red disk. Thus, in general, $k=\Delta/2$. In the arrangement, we can identify $2k-1=\Delta-1$ rings, each consisting of a red and a blue face at equal distance from the innermost blue face; see the right image in Fig. 4, where the rings are shown in black and gray. Going from the innermost blue face to the innermost ring requires two hops, going from a ring to the next ring requires two hops, and going from the outermost ring to the outermost blue face requires two hops. Thus, the distance between the innermost and outermost blue face is $2\Delta$.
This construction gives an arrangement of two disks with overlap number $\Delta>1$ whose dual graph has diameter $2\Delta$, for even $\Delta\geqslant 2$. By omitting the last two corridors of both disks, we obtain such an arrangement for odd $\Delta>2$ as well.
3 The general case of $n$ disks
We now consider the case of general $n$, which turns out to be significantly harder than the case of two disks. The main difficulty lies in bounding the number of maximal faces; for two disks this number is simply the overlap number $\Delta$ (assuming the disks are not disjoint), but for general $n$ it is not even clear that the number of maximal faces (or the number of maximum faces, for that matter) is bounded. Our main effort lies in proving that this is the case (Section 3.1). Once we have shown that, establishing a bound on the diameter of the dual graph of the arrangement of disks is relatively easy (Section 3.2).
3.1 The proof of Theorems 1.2 and 1.3
Recall that $\mu(A)$ and $M(\mathcal{A})$ denote the number of maximal and maximum faces, respectively, in the arrangement $\mathcal{A}$ defined by the set $\mathcal{D}$ of $n$ disks. We first prove our bound on the number of maximum faces. See 1.2
Proof 3.1.
Clearly, $M(2,\Delta)=\Delta$ by definition. We will show that for $n\geqslant 3$ we have
$M(n,\Delta)\leqslant M(n-1,\Delta)+4(n-1)(\Delta-1).$ (1)
This implies the theorem, since
$\begin{array}[]{lll}M(n,\Delta)&\leqslant&M(2,\Delta)+\sum_{j=3}^{n}4(j-1)(\Delta-1)\\ &=&\Delta+\sum\limits_{j=3}^{n}4(j-1)(\Delta-1)\\ &<&\sum\limits_{j=2}^{n}4(j-1)\Delta\mbox{ }\\ &=&2n(n-1)\Delta.\end{array}$
To prove that Inequality (1) holds, consider a set $\mathcal{D}$ of $n\geqslant 3$ disks with overlap number $\Delta$, and let $D_{0}$ be an arbitrary disk in $\mathcal{D}$. Define $\mathcal{D}_{1}:=\mathcal{D}\setminus\{D_{0}\}$ and let $D_{1},\ldots,D_{n-1}$ be the disks in $\mathcal{D}_{1}$. We will show that
$M(\mathcal{A}(\mathcal{D}))\leqslant M(\mathcal{A}(\mathcal{D}_{1}))+4(n-1)(\Delta-1).$
Note that some faces do not change when we remove $D_{0}$ from $\mathcal{D}$, while other faces—the ones incident to $\partial D_{0}$—expand into larger faces. Now consider a maximum face $f$ of $\mathcal{A}(\mathcal{D})$ and let $\mathit{parent}(f)$ be the face in $\mathcal{A}(\mathcal{D}_{1})$ that contains $f$. Note that $\mathit{parent}(f)$ is a maximum face in $\mathcal{A}(\mathcal{D}_{1})$. We consider two types of maximum faces in $\mathcal{A}(\mathcal{D})$, illustrated in Figure 5(i).
Figure 5: (i) An arrangement of three disks. Faces $f_{1}$ and $f_{4}$ are safe, faces $f_{2}$ and $f_{3}$ are unsafe. (ii) The red disk defines four dangerous intervals along $\partial D_{0}$, indicated by thick lines.
  • A safe face is a maximum face $f$ in $\mathcal{A}(\mathcal{D})$ such that $f$ is the unique maximum face in $\mathcal{A}(\mathcal{D})$ contained in $\mathit{parent}(f)$.
  • An unsafe face is a maximum face $f$ in $\mathcal{A}(\mathcal{D})$ such that there is a maximum face $f^{\prime}\neq f$ in $\mathcal{A}(\mathcal{D})$ that is also contained in $\mathit{parent}(f)$.
The number of safe faces is obviously at most $M(\mathcal{D}_{1})$, so
[…]
Note that the inequality is strict, unless there are no unsafe faces. Indeed, any set of unsafe faces with the same parent face already contributes 1 to $M(\mathcal{D}_{1})$. Next, we bound the number of unsafe faces.
Consider the arrangement $\mathcal{A}(\{D_{0},D_{i}\})$ induced by $D_{0}$ and $D_{i}$. Any vertex $v$ of this arrangement is incident to two edges of $\mathcal{A}(\{D_{0},D_{i}\})$ contained in $\partial D_{i}$, one inside $D_{0}$ and one outside $D_{0}$. We call the latter edge the outside curve of $v$ and we denote it by $\gamma(v)$. Let $\mathcal{I}(D_{i}):=D_{i}\cap\partial D_{0}$ be the (possibly empty) set of intervals on $\partial D_{0}$ covered by $D_{i}$. Consider an interval $I\in\mathcal{I}(D_{i})$ and let $v$ and $w$ be its endpoints. We call the interval $I$ dangerous if $\gamma(v)\neq\gamma(w)$; see Figure 5(ii) for an example.
Claim 1.
Any disk $D_{i}\in\mathcal{D}_{1}$ defines at most $2\Delta-2$ dangerous intervals.
{claimproof}
Assume there is at least one dangerous interval, otherwise the claim holds trivially. Let $F_{i}$ be the set of faces of $\mathcal{A}(\{D_{0},D_{i}\})$ that are incident to a dangerous interval, let $F_{i}^{\mathrm{in}}\subset F_{i}$ be the faces that lie inside $D_{0}$ and let $F_{i}^{\mathrm{out}}\subset F_{i}$ be the faces that lie outside $D_{0}$. Let $\mathcal{G}$ be the graph whose nodes are the faces in $F_{i}$ and where there is an edge between two faces if they share a dangerous interval; see Figure 6.
Figure 6: The graph $\mathcal{G}$ used in the proof of 1.
$\mathcal{G}$ is a bipartite graph with node sets $F_{i}^{\mathrm{in}}$ and $F_{i}^{\mathrm{out}}$, and that $|F_{i}^{\mathrm{in}}|\leqslant\Delta$ by definition of the overlap number. Moreover, any node $f\in F_{i}^{\mathrm{out}}$ has degree at least two in $\mathcal{G}$. Indeed, if $f$ had degree one, then the face $f$ would be adjacent to a single dangerous interval, which is impossible by the definition of dangerous intervals. Note that $\mathcal{G}$ is a tree, since $D_{i}$ is a topological disk.
We now view $\mathcal{G}$ as a rooted tree, where we pick an arbitrary node $f\in F_{i}^{\mathrm{out}}$ as the root. Since the nodes in $F_{i}^{\mathrm{out}}$ have degree at least 2 in $\mathcal{G}$, they are not leaves. To bound $|F_{i}^{\mathrm{out}}|$, we charge each node in $F_{i}^{\mathrm{out}}$ to an arbitrary one of its children. Thus we charge each node in $F_{i}^{\mathrm{out}}$ to a node in $F_{i}^{\mathrm{in}}$ in such a way that each node in $F_{i}^{\mathrm{in}}$ is charged at most once. Since the root node has at least two children, one of which will not be charged, we conclude that
$|F_{i}^{\mathrm{out}}|\leqslant|F_{i}^{\mathrm{in}}|-1\leqslant\Delta-1.$
Hence, the total number of nodes in $\mathcal{G}$ is at most $2\Delta-1$, and so the number of edges in $\mathcal{G}$, which equals the number of dangerous intervals, is at most $2\Delta-2$. We call an edge of a face $f$ of $\mathcal{A}(\mathcal{D})$ a boundary edge of $f$ if it is contained in $\partial D_{0}$. The next claim relates the number of dangerous intervals to the number of unsafe faces.
Claim 2.
Let $f$ be an unsafe face. Then $f$ has a boundary edge $e$ such that at least one of the endpoints of $e$ is an endpoint of a dangerous interval.
{claimproof}
Let $f^{\prime}\neq f$ be a maximum face in $\mathcal{A}(\mathcal{D})$ that is contained in $\mathit{parent}(f)$; such a face exists because $f$ is unsafe. Then there is a curve $\xi$ in the interior of $\mathit{parent}(f)$ that connects $f$ to $f^{\prime}$. (More precisely, $\xi$ connects a point in the interior of $f$ to a point in the interior of $f^{\prime}$.) Since $f$ and $f^{\prime}$ are distinct faces in $\mathcal{A}(\mathcal{D})$, this curve $\xi$ must cross a boundary edge, $e$, of $f$. We can assume without loss of generality that $\xi$ crosses $e$ once. Let $v$ and $w$ be the endpoints of $e$, and assume for a contradiction that neither $v$ nor $w$ is an endpoint of a dangerous interval. Let $\gamma(v)$ and $\gamma(w)$ be the outside curves of $v$ and $w$, respectively, and let $D_{i}$ and $D_{j}$ be such that $\gamma(v)\subset\partial D_{i}$ and $\gamma(w)\subset\partial D_{j}$ We consider two cases.
Figure 7: Illustration for the proof of 2.
  • If $i=j$ then $\gamma(v)=\gamma(w)$, because we assumed that neither $v$ nor $w$ is an endpoint of a dangerous interval. Hence, $f^{\prime}$ is contained in the region $R$ enclosed by $\gamma(v)\cup e$; see Figure 7(i). But $R$ lies fully outside $D_{0}$, which contradicts that $f^{\prime}$ is a maximum face in $\mathcal{A}(\mathcal{D})$.
  • If $i\neq j$ then $\gamma(v)$ must intersect $\gamma(w)$. To see this, consider the interval $I_{v}\subset\partial D_{0}\cap D_{i}$ with endpoint $v$ and containing $e$, and the interval $I_{w}\subset\partial D_{0}\cap D_{j}$ with endpoint $w$ and containing $e$. Then $w\in I_{v}$ and $v\in I_{w}$. Together with the fact that $I_{v}$ and $I_{w}$ are not dangerous by assumption, this implies that $\gamma(v)$ indeed intersects $\gamma(w)$; see Figure 7(ii). Let $R$ be the bounded face in the arrangement defined by $e$, $\gamma(v)$ and $\gamma(w)$ that is incident to $e$. Then $f^{\prime}$ is contained in $R$, which lies outside $D_{0}$, contradicting that $f^{\prime}$ is a maximum face in $\mathcal{A}(\mathcal{D})$.
We derived a contradiction in both cases, thus finishing the proof of the claim. 2 allows us to charge each unsafe face to a dangerous interval in such a way that each unsafe interval is charged at most once for each of its endpoints. By 1, the total number of dangerous intervals is at most $(n-1)(2\Delta-2)$. Hence, the number of unsafe faces is at most $2(n-1)(2\Delta-2)$. This finishes the proof of Inequality (1) and, hence, of the theorem.
The construction in Figure 8 shows that $M(n,\Delta)=\Omega(n^{2}\Delta)$.
Figure 8: A construction showing that $M(n,\Delta)=\Omega(n^{2}\Delta)$.
Theorem 1.2 implies the following bound on the number of maximal faces.
See 1.3
Proof 3.2.
We claim that for any $n\geqslant 2$ we have $\mu(n,\Delta)\leqslant(2^{n}-1)\cdot M(n,\Delta)$. Together with Theorem 1.2, this implies the theorem.
To prove the claim, let $\mathcal{D}$ be a collection of $n$ topological disks with overlap number $\Delta$, and let $f$ be a maximal face of $\mathcal{A}(\mathcal{D})$. Then $f$ is a maximum face in the arrangement $\mathcal{A}(\mathcal{D}_{f})$, where $\mathcal{D}_{f}\subseteq\mathcal{D}$ is the set of disks containing $f$. There are at most $M(|\mathcal{D}_{f}|,\Delta)$ such faces, which we charge to $D_{f}$. Hence,
$\mu(n,\Delta)\leqslant\sum_{\emptyset\neq\mathcal{D}^{\prime}\subseteq\mathcal{D}}M(|\mathcal{D}^{\prime}|,\Delta)\leqslant\sum_{\emptyset\neq\mathcal{D}^{\prime}\subseteq\mathcal{D}}M(n,\Delta)=(2^{n}-1)\cdot M(n,\Delta),$
as claimed. For the second inequality, note that $M(n,\Delta)$ is non-decreasing in $n$.
3.2 The proof of Theorem 1.4
Using the bound on the number of maximal faces in the arrangement $\mathcal{A}$, we now obtain a bound on the diameter of the dual graph $\mathcal{G}^{*}$ of $\mathcal{A}$.
Recall that the ply of a face $f$ in the arrangement $\mathcal{A}$, denoted by $\operatorname{ply}(f)$, is the total number of disks in $\mathcal{D}$ that contain $f$. Define $p_{\max}:=\max_{f}\operatorname{ply}(f)$ to be the maximum ply over all faces in the arrangement $\mathcal{A}$. For a node $u$ in the dual graph $\mathcal{G}^{*}$, we define $\operatorname{ply}(u):=\operatorname{ply}(f_{u})$, where $f_{u}$ is the face in $\mathcal{A}$ corresponding to $u$. The next lemma relates $\mu(\mathcal{A})$, the number of maximal faces in $\mathcal{A}$, to $\operatorname{diam}(\mathcal{G}^{*})$, the diameter of the dual graph $\mathcal{G}^{*}$ of $\mathcal{A}$. Combining the lemma with Theorem 1.3 proves Theorem 1.4.
Lemma 3.3.
Let $\mathcal{A}$ be an arrangement induced by a set $\mathcal{D}$ of $n$ disks, and let $\mathcal{G}^{*}$ be the dual graph of $\mathcal{A}$. Then $\operatorname{diam}(\mathcal{G}^{*})\leqslant 2p_{\max}\cdot\mu(\mathcal{A})\leqslant 2n\cdot\mu(\mathcal{A})$.
Proof 3.4.
Let $V$ be the set of nodes of $\mathcal{G}^{*}$, and let $M=\{v_{1},\ldots,v_{t}\}$ be the subset of nodes from $V$ corresponding to maximal faces in $\mathcal{A}$. We say that a node $u\in V$ has a monotone path to a node $v_{i}\in M$ if there is a path $u=w_{1},\ldots,w_{s}=v_{i}$ in $\mathcal{G}^{*}$ such that $\operatorname{ply}(w_{j+1})=\operatorname{ply}(w_{j})+1$ for all $1\leqslant j<s$. Observe that a node $v_{i}\in M$ has a monotone path to itself and to no other node $v_{j}\in M$. We partition $V$ into subsets $V_{1},\ldots,V_{t}$, one for each node $v_{i}\in M$, where we define
[…]
Note that each node $u\in V$ is assigned to a set $V_{i}$, because from $u$ we can greedily move to faces of higher ply until we reach a maximal face. Also note that $v_{i}\in V_{i}$ for all $i$.
Now, let $\overline{\mathcal{G}^{*}}$ be the graph obtained from $\mathcal{G}^{*}$ by contracting each subgraph $\mathcal{G}^{*}[V_{i}]$ to a single node. We make the following two observations.
  1. (i)
    $\overline{\mathcal{G}^{*}}$ is a connected graph (because $\mathcal{G}^{*}$ is obviously connected) with $\mu(\mathcal{A})$ nodes. Hence, the distance between any two nodes $V_{i},V_{j}$ in $\overline{\mathcal{G}^{*}}$ is at most $\mu(\mathcal{A})-1$.
  2. (ii)
    Let $v_{i},v_{j}\in M$ be such that the nodes in $\overline{\mathcal{G}^{*}}$ corresponding to the sets $V_{i}$ and $V_{j}$ are adjacent. Then, the distance $d(v_{i},v_{j})$ between the nodes $v_{i}$ and $v_{j}$ in $\mathcal{G}^{*}$ is at most $\operatorname{ply}(v_{i})+\operatorname{ply}(v_{j})$. To see this, let $w_{i}\in V_{i}$ and $w_{j}\in V_{j}$ be adjacent nodes in $\mathcal{G}^{*}$. Such nodes must exist because $V_{i}$ and $V_{j}$ are adjacent in $\overline{\mathcal{G}^{*}}$. From $w_{i}$ we can walk to $v_{i}$ in $\operatorname{ply}(v_{i})-\operatorname{ply}(w_{i})$ hops and from $w_{j}$ we can walk to $v_{j}$ in $\operatorname{ply}(v_{j})-\operatorname{ply}(w_{j})$ hops. Hence,
    $d(v_{i},v_{j})\leqslant\big(\operatorname{ply}(v_{i})-\operatorname{ply}(w_{i})\big)+\big(\operatorname{ply}(v_{j})-\operatorname{ply}(w_{j})\big)+1\leqslant\operatorname{ply}(v_{i})+\operatorname{ply}(v_{j}),$
    where the last inequality follows because $\operatorname{ply}(w_{i})$ and $\operatorname{ply}(w_{j})$ cannot both be zero.
Now, let $u,u^{\prime}$ be two arbitrary nodes in $V$. Let $i,j$ be such that $u\in V_{i}$ and $u^{\prime}\in V_{j}$. We can walk from $u$ to $v_{i}$ in $\operatorname{ply}(v_{i})-\operatorname{ply}(u)\leqslant p_{\max}$ hops, and we can walk from $u^{\prime}$ to $v_{j}$ in $\operatorname{ply}(v_{j})-\operatorname{ply}(u^{\prime})\leqslant p_{\max}$ hops. Moreover, by combining the two observations above, we see that there is a path from $v_{i}$ to $v_{j}$ in $\mathcal{G}^{*}$ consisting of $(\mu(\mathcal{A})-1)\cdot 2p_{\max}$ edges. Hence, $d(u,u^{\prime})\leqslant 2p_{\max}\cdot\mu(\mathcal{A})$. It remains to observe that $p_{\max}\leqslant n$, which we trivially have.
4 Concluding Remarks and Discussion
In this paper, we derived bounds on the diameter of the dual graph of an arrangement $\mathcal{A}$ of disks in terms of the number of disks, $n$, and the maximum number of intersection components between any two disks, $\Delta$. (The bound on the diameter is equivalent to a bound on the number of crossings of any minimum-crossing $st$-curve in arrangements of two disks.)
For an arrangement containing only two disks, we established a sharp linear bound of $\max\{2,2\Delta\}$. For an arrangement of $n$ disks, we bounded the diameter of the dual graph by $O(n^{3}2^{n}\Delta)$. This was done by establishing an $O(n^{2}2^{n}\Delta)$ bound on the number of maximal faces in $\mathcal{A}$. The latter was done by first proving an $O(n^{2}\Delta)$ bound on the number of maximum faces, which is asymptotically tight.
Our bound on the number of maximal faces is far from tight: our best lower bound is the same as for the number of maximum faces, namely $\Omega(n^{2}\Delta)$, while our upper bound is $O(n^{2}2^{n}\Delta)$. The main open problem is to tighten this gap. We believe that the true bound is polynomial in nn—in fact, we conjecture that the $\Omega(n^{2}\Delta)$ lower bound is asymptotically tight. The upper bound on the number of maximal faces was used to derive a bound on the diameter of the dual graph of the arrangement. Notice that by making $n/2$ nested copies of each of the two disks in Figure 4, we can produce an arrangement of $n$ disks with overlap $\Delta$ and diameter $\Omega(n\Delta)$. Instead of trying to improve the $O(n^{3}2^{n}\Delta)$ upper bound by improving the bound on the number of maximal faces, one may also try to prove a better bound directly.